A lift and its passengers with a total mass of 500 kg accelerates upwards at 2ms-2 as shown.

Assume that g = 10 ms-2.


1
Expert's answer
2021-01-29T04:29:40-0500

The tension in the cable that makes the elevator move can be found by Newton's second law:


"ma=T-mg,\\
T=m(a+g)=500\cdot12=6000\text{ N}."